NNCuts on LRHWW, Mll and MetSig

Richard St. Denis University of Glasgow/CDF 17 December 2007

H → WW Meeting

[figure]

Outline

▶ MetSig, Met and SumEt by jet and Hi/Lo S/B ▶ BaseSameSign

H → WW Meeting

[figure: SumEt]

Richard St. Denis, Glasgow University/CDF, 10 December 2

H → WW Meeting

[figure: SumEt]

Richard St. Denis, Glasgow University/CDF, 10 December 3

H → WW Meeting

[figure: DiMass]

Richard St. Denis, Glasgow University/CDF, 10 December 4

H → WW Meeting

[figure: DiMass]

Richard St. Denis, Glasgow University/CDF, 10 December 5

H → WW Meeting

[figure: Met]

Richard St. Denis, Glasgow University/CDF, 10 December 6

H → WW Meeting

[figure: Met]

Richard St. Denis, Glasgow University/CDF, 10 December 7

H → WW Meeting

[figure: MetSig]

Richard St. Denis, Glasgow University/CDF, 10 December 8

H → WW Meeting

[figure: MetSig]

Richard St. Denis, Glasgow University/CDF, 10 December 9

H → WW Meeting

[figure]

dPhiLeptons

Richard St. Denis, Glasgow University/CDF, 10 December 10

H → WW Meeting

[figure]

dPhiLeptons

Richard St. Denis, Glasgow University/CDF, 10 December 11

H → WW Meeting

[figure]

dRLeptons

Richard St. Denis, Glasgow University/CDF, 10 December 12

H → WW Meeting

[figure]

dRLeptons

Richard St. Denis, Glasgow University/CDF, 10 December 13

H → WW Meeting

[figure: jet1 Et]

Richard St. Denis, Glasgow University/CDF, 10 December 14

H → WW Meeting

[figure: jet1 Et]

Richard St. Denis, Glasgow University/CDF, 10 December 15

H → WW Meeting

[figure]

Base SameSign 0 jet

Richard St. Denis, Glasgow University/CDF, 10 December 16